OpenStudy (christos):

can you find the derivative of y = ln|1 + x| ? Is there a caught when we have somehing like |....| ?

OpenStudy (anonymous):

\[\Large\frac{d }{dx} \ln f(x)=\frac{ f'(x) }{ f(x) }\]

OpenStudy (anonymous):

there |...| is unneccessary since the defination of ln functi0on desires it to be positive

OpenStudy (anonymous):

(lnabs(1+x))=(ln(x+1) )'=1/(x+1)

OpenStudy (anonymous):

What is the anti-derivative of 1/x?

OpenStudy (anonymous):

lnabsx

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

Then if I differentiate that, I get back to 1/x. Similarly, what would be the antiderivative of 1/x+1? What would be the derivative of this antiderivative that you just got?

OpenStudy (anonymous):

@Christos

OpenStudy (raden):

y = |1+x| y = sqrt((1+x)^2) squares to both sides : y^2 = (1+x)^2 now, take the derivative by using implicit, 2y dy = 2(1+x)(1)dx dy/dx = 2(1+x)/2y dy/dx = (1+x)/y subtitute back, that y = |1+x| so, dy/dx = (1+x)/|1+x|

OpenStudy (raden):

does that make sense ?

OpenStudy (christos):

thank you